What's the x-intercept and y-intercept of the equation 2y+x= 12
Found 2 solutions by Cromlix, Yanisleysi :
Answer by Cromlix(4381)

Sort equation into y = mx + c
2y + x = 12 => y = -1/2 x + 6
This line cuts the y axis when x = 0
y = -1/2(0) + 6 => (0,6)
It cuts the x axis when y = 0
0 = -1/2x + 6
-1/2x = -6
1/2x = 6
x = 12 => (12,0)
Hope this helps
:-)

Answer by Yanisleysi (5)

2y+x= 12
2y=x-12